### v1

# Definable henselian valuations in positive residue characteristic

Abstract.
We study the question of $\mathcal{L}_{\mathrm{ring}}$-definability of non-trivial henselian valuation rings. Building on previous work of Jahnke and Koenigsmann, we provide a characterization of henselian fields that admit a non-trivial definable henselian valuation. In particular, we treat cases where the canonical henselian valuation has positive residue characteristic, using techniques from the model theory and algebra of tame fields.
1. Introduction
Many successful investigations around the existence and properties of solutions of polynomial equations over a field rely on enriching the field with a valuation. For example over the field of $p$-adic numbers $\mathds{Q}_{p}$, the valuation ring can actually be completely recovered as an $\mathcal{L}_{\mathrm{ring}}$-definable subset, as first observed in the seminal work of Julia Robinson on Hilbert’s Tenth Problem [Rob63]. Understanding this phenomenon is a classical topic of the model theory of valued fields, and it has striking applications in the investigations around dividing lines for fields, see e.g. Johnson’s spectacular classification of dp-finite fields in [Joh20]. For a more thorough survey on definability of valuations we refer the reader to [FJ17].
In this manuscript, we will focus on the specific problem of characterizing which fields admit a non-trivial definable henselian valuation. There are some clear obstructions: for example, if the field is separably closed, then the answer is always no. Once this case is excluded, Jahnke and Koenigsmann isolate in [JK17] necessary and sufficient properties of the canonical henselian valuation $v_{K}$ of a field $K$ (see Subsection 2.3) for the existence of a definable non-trivial henselian valuation, under the assumption that $\operatorname{char}{Kv_{K}}=0$.
Theorem 1.0.1 ([JK17, Corollary 6.1]).
Let $K$ be a henselian field that is not separably closed, with $\operatorname{char}{Kv_{K}}=0$. Then $K$ admits a definable non-trivial henselian valuation if and only if at least one of the following conditions hold:
  1. (1)
    $Kv_{K}$ is separably closed,
  2. (2)
    $Kv_{K}$ is not t-henselian,
  3. (3)
    there is $L\succeq Kv_{K}$ such that $v_{L}L$ is not divisible,
  4. (4)
    $v_{K}K$ is not divisible.
This characterization (and its extension, which we will state in a moment) falls into the general philosophy of Ax-Kochen/Ershov principles. Typically, (model-theoretic) questions about a valued field $(K,v)$ are answered using its residue field $Kv$ and its value group $vK$. In this case, since the question is about a field $K$ with no specified valuation, one ought to isolate a canonical one among all its henselian valuations; this is usually denoted by $v_{K}$ (see Subsection 2.3). Answers to questions about $K$, then, should be given in terms of properties of $Kv_{K}$ and $v_{K}K$.
1.1. Our result
We extend their result to remove the assumption on the residue characteristic. We have to restrict ourselves to either perfect positive characteristic fields, or fields of characteristic zero for which $\mathcal{O}_{v_{K}}/p$ is semi-perfect. While the result might be sharpened by weakening the assumption in mixed characteristic, an understanding of imperfect valued fields of positive characteristic seems to be currently out of reach (except for few examples, e.g. the separably tame ones).
Main theorem .
Let $K$ be perfect, not separably closed, and henselian. If ${\operatorname{char}{K}=0}$ and ${\operatorname{char}{Kv_{K}}=p>0}$, assume that $\mathcal{O}_{v_{K}}/p$ is semi-perfect. Then $K$ admits a definable non-trivial henselian valuation if and only if at least one of the following conditions hold:
  1. (1)
    $Kv_{K}$ is separably closed,
  2. (2)
    $Kv_{K}$ is not t-henselian,
  3. (3)
    there is $L\succeq Kv_{K}$ such that $v_{L}L$ is not divisible,
  4. (4)
    $v_{K}K$ is not divisible,
  5. (5)
    $(K,v_{K})$ is not defectless,
  6. (6)
    there is $L\succeq Kv_{K}$ such that $(L,v_{L})$ is not defectless.
Remark 1.1.1 ().
Note that in conditions 3 and 6, one automatically has that $v_{L}$ is non-trivial (since the trivial valuation has divisible value group, and is defectless).
In the proof we draw in one way or another from the technology of defect and tame valued fields. The paper is structured as follows:
  • $\circ$
    Section 2 sets the stage, by fixing the notation and a few elementary facts.
  • $\circ$
    Section 3 explores the two roles played by divisibility of the value group: on the one hand, in Proposition 3.2.1, failure of divisibility in an elementary extension of $Kv_{K}$ is exploited to ($\varnothing$-)define a valuation; on the other, in Theorem 3.1.5, divisibility is the keystone to finding obstructions to the existence of definable valuations.
  • $\circ$
    Section 4 deploys the technology of independent defect to build definable henselian valuations from certain kinds of Galois defect extensions.
  • $\circ$
    Section 5 puts everything together, providing the full characterization as advertised.
1.2. Acknowledgements
This work was done as part of the PhD projects of the authors: we therefore wish to thank Sylvy Anscombe, Arno Fehm, Martin Hils, and Franziska Jahnke for their perfect guidance, highly valued encouragement, and dedication in finding and overcoming the defects in our arguments. We would like to express special thanks to S. Anscombe for pointing us towards Beth definability in time of need, to F. Jahnke for directing us towards and through these questions, and suggesting the use of independent defect, and to A. Fehm for pointing out how the proof of Proposition 3.2.1 could be simplified and applied in a more general setting. We also wish to thank Blaise Boissonneau, Nicolas Daans, Akash Hossain, and Franz-Viktor Kuhlmann for fruitful discussions.
This project began when the third author was a Visiting Doctoral Researcher to the Cluster of Excellence “Mathematics Münster”. This research is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under the Excellence Strategy EXC 2044–390685587, Mathematics Münster: Dynamics–Geometry–Structure. The first two authors are also funded by the ANR-DFG program AAPG2019, GeoMod: Geometric and combinatorial configurations in Model theory. The second author is also funded by the DFG project 495759320 – Modelltheorie bewerteter Körper mit Endomorphismus. The third author is also funded by the DFG project 404427454 – Definierbarkeit und Entscheidbarkeit in globalen und lokalen Körpern.
2. Preliminaries
2.1. Notation
Given a valued field $(K,v)$, we denote by $\mathcal{O}_{v}$ or $\mathcal{O}_{(K,v)}$ its valuation ring with maximal ideal $\mathfrak{m}_{v}$, $Kv$ its residue field and $vK$ its value group. We will write ${vK}_{\infty}$ for $vK\cup\{\infty\}$. We will denote by $\operatorname{res}_{v}\colon\mathcal{O}_{v}\mathchoice{\longrightarrow}{\rightarrow}{\rightarrow}{\rightarrow}Kv$ the residue map. Given a field $K$, we denote its separable closure by $K^{\mathrm{sep}}$. A field will be called henselian if it admits a non-trivial henselian valuation. A ring of characteristic $p>0$ is called semi-perfect if $x\mapsto x^{p}$ is surjective (but not necessarily injective). If $K\subseteq L$ is a Galois extension, we will write $\operatorname{Gal}(L|K)$ for the corresponding Galois group. We will mostly work with the two languages $\mathcal{L}_{\mathrm{ring}}=\{+,\;\cdot\;,0,1\}$, and $\mathcal{L}_{\mathrm{val}}=\mathcal{L}_{\mathrm{ring}}\cup\{\mathcal{O}\}$, where $\mathcal{O}$ is a unary predicate. If $M$ is a first-order structure, $I$ is an index set and $\mathcal{U}$ is a ultrafilter on $I$, then we denote by $M^{\mathcal{U}}$ the corresponding ultrapower. Unless otherwise stated, definable will mean definable with parameters. Otherwise, we say $\varnothing$-definable.
2.2. Coarsenings and compositions
We briefly recall some facts about coarsenings and refinements of valuations, which are explained in greater depth in [EP05, Section 2.3]. For two valuations $v$ and $w$ on $K$, we say that $v$ is finer than $w$ or is a refinement of $w$ (that $w$ is coarser than $v$ or is a coarsening of $v$), if $\mathcal{O}_{v}\subseteq\mathcal{O}_{w}$. We identify $v$ and $w$ if they have the same valuation ring, and use $v$ and $\mathcal{O}_{v}$ interchangeably. We say $v$ and $w$ are comparable if $v$ is a coarsening of $w$, or $w$ is a coarsening of $v$. Coarsenings of a fixed valuation ring $\mathcal{O}_{v}$ are linearly ordered by inclusion and in one-to-one correspondence with convex subgroups of $vK$. We denote by $v_{H}$ the coarsening of $v$ associated to a convex subgroup $H$ of $vK$.
Given a valuation $v$ and a coarsening $w$ corresponding to the convex subgroup $H\subseteq vK$, we denote the induced valuation $\bar{v}\colon(Kw)^{\times}\mathchoice{\longrightarrow}{\rightarrow}{\rightarrow}{\rightarrow}H$ on $Kw$ by $\bar{v}$, with value group $H$ and residue field $Kv$. Given a valuation $w$ on $K$ and a valuation $u$ on its residue field $Kw$, we define the composition $v=u\circ w$ to be the valuation with valuation ring $\mathcal{O}_{v}=\operatorname{res}_{w}^{-1}(\mathcal{O}_{u})$. It has residue field $Kv=(Kw)u$ and value group $u(Kw)$, a convex subgroup of $vK$ with $vK/u(Kw)\cong wK$.
We can write the valuations as places and obtain
{tikzcd}
2.3. The canonical henselian valuation
What follows is classical and can be found in greater detail in [EP05, Section 4.4]. For any field $K$, one can arrange (equivalence classes of) henselian valuations on $K$ according to whether their residue field is separably closed or not. Namely, one can define
$H_{1}(K)=\{v\mid(K,v)\text{ henselian and }(Kv)^{\mathrm{sep}}\neq Kv\}$
and
$H_{2}(K)=\{v\mid(K,v)\text{ henselian and }(Kv)^{\mathrm{sep}}=Kv\}.$
As above, we identify $v$ with $\mathcal{O}_{v}$, so we will often write $\mathcal{O}_{v}\in H_{1}(K)$ to mean $v\in H_{1}(K)$.
The set $H_{1}(K)$ is linearly ordered by coarsening. If $H_{2}(K)=\varnothing$, then there is a finest valuation in $H_{1}(K)$, which we denote by $v_{K}$. Otherwise, we let $v_{K}$ be the coarsest valuation in $H_{2}(K)$. We call $v_{K}$ the canonical henselian valuation on $K$.
{tikzpicture}
Every henselian valuation on $K$ is comparable to $v_{K}$. Moreover, $v_{K}$ is non-trivial if and only if $K$ is henselian and not separably closed. By definition, ${Kv_{K}\neq(Kv_{K})^{\mathrm{sep}}}$ if and only if $v_{K}\in H_{1}(K)$ if and only if $H_{2}(K)=\varnothing$, and if these equivalent conditions hold, then all henselian valuation rings on $K$ are linearly ordered and coarser than $\mathcal{O}_{v_{K}}$.
Remark 2.3.1 ().
Let $(K,v)$ be a henselian valued field with $v\in H_{1}(K)$. Then, ${v_{K}=v_{Kv}\circ v}$. Indeed, there is a correspondence between the set $H$ of henselian refinements of $v$ on $K$ and valuations on $Kv$. It is given by sending any refinement $u$ of $v$ to the induced valuation $\overline{u}$ on $Kv$, as in the following diagram (where valuations are written as places):
[…]
Since the correspondence preserves residue fields, $H\cap H_{1}(K)$ is mapped to $H_{1}(Kv)$, and similarly for $H\cap H_{2}(K)$.
{tikzpicture}
The correspondence preserves the coarsening relation. In particular, it follows that $v_{K}$ is mapped to $v_{Kv}$ and $v_{K}=v_{Kv}\circ v$.
2.4. Defect and tame fields
Let $(K,v)\subseteq(L,w)$ be a finite extension of a henselian valued field. Let $p$ be the characteristic exponent of $Kv$, namely $p=1$ if $\operatorname{char}(Kv)=0$, and $p=\operatorname{char}(Kv)$ otherwise. Then one has, by [EP05, Theorem 3.3.3],
($\star$) $[L:K]=d\cdot(wL:vK)\cdot[Lw:Kv],$
for some $d=p^{\nu}\in\mathds{N}$. We say that the extension is defectless if $d=1$ (or equivalently $\nu=0$), otherwise we say that the extension has defect. The valued field $(K,v)$ is called defectless if all of its finite extensions are defectless. We refer to [Kuh16] for the details on the various equivalent definitions of “tame”.
Definition 2.4.1 ().
Let $(K,v)$ be a henselian valued field and let $p$ be the characteristic exponent of the residue field. Then $(K,v)$ is called tame if it is algebraically maximal, $vK$ is $p$-divisible, and $Kv$ is perfect. Equivalently, $(K,v)$ is tame if it is defectless, $vK$ is $p$-divisible, and $Kv$ is perfect.
In the case of valued fields where $p=1$ (i.e., valued fields of equicharacteristic zero), tame is equivalent to henselian.
Remark 2.4.2 (see [Kuh16, Section 7]).
The class of tame valued fields is an elementary class in the language $\mathcal{L}_{\mathrm{val}}$.
2.5. t-henselianity and saturation
A field $K$ is said to be t-henselian if it is elementarily equivalent, in $\mathcal{L}_{\mathrm{ring}}$, to a field $L$ admitting a non-trivial henselian valuation. However, a field $K$ can be t-henselian without being henselian. The next fact is well-known, and its proof follows from [PZ78, Lemma 3.3] and [Pre91, Page 203].
Fact 2.5.1 ().
Let $K$ be t-henselian and $\aleph_{1}$-saturated. Then $K$ admits a non-trivial henselian valuation. In particular, if $K$ is t-henselian and $K\preceq L$ is an $\aleph_{1}$-saturated elementary extension, then $L$ admits a non-trivial henselian valuation.
2.6. Definability in Jahnke-Koenigsmann
For use in later sections, we also summarize two of the main theorems of [JK17], which will prove to be fundamental tools in our work. In particular, we highlight how the second part of the upcoming statement allows us to often assume that the value group we are working with is divisible.
Theorem 2.6.1 ([JK17, Theorems A and B]).
Let $K$ be a henselian field which is not separably closed. Assume that $K$ satisfies at least one of the following:
  1. (1)
    $Kv_{K}$ is separably closed,
  2. (2)
    $Kv_{K}$ is not t-henselian,
  3. (3)
    there is a prime $p$ such that $v_{K}K$ is not $p$-divisible and not $p$-antiregular,
then $K$ admits a $\varnothing$-definable non-trivial henselian valuation. In general, if $v_{K}K$ is not divisible, then $K$ admits a definable (possibly with parameters) non-trivial henselian valuation.
3. Divisibility
3.1. Divisibility as an obstruction
The proof of Theorem 1.0.1 uses the induced structure on the value group of an equicharacteristic zero henselian valued field to argue against definability of valuations. We use the work of Jahnke and Simon in [JS20] to generalize this argument.
Definition 3.1.1 ().
Let $\mathcal{L}_{{K,\Gamma,k}}$ be a three-sorted language with sorts $K$ (with language $\mathcal{L}_{\mathrm{ring}}$), $\Gamma$ (with language $\mathcal{L}_{\mathrm{oag}}\cup\{\infty\}$) and $k$ (with language $\mathcal{L}_{\mathrm{ring}}$), and functions $v\colon K\mathchoice{\longrightarrow}{\rightarrow}{\rightarrow}{\rightarrow}\Gamma$ and $\operatorname{res}\colon K\mathchoice{\longrightarrow}{\rightarrow}{\rightarrow}{\rightarrow}k$. Any valued field $(K,v)$ can be interpreted as an $\mathcal{L}_{{K,\Gamma,k}}$-structure in the natural way.
Fact 3.1.2 ([JS20, Lemma 3.1]).
Let $(K,v)$ be a separably tame field of positive residue characteristic and finite degree of imperfection. Then the value group $vK$ is purely stably embedded as an ordered abelian group, i.e. for every $D\subseteq vK^{n}$ which is $\mathcal{L}_{{K,\Gamma,k}}$-definable (possibly with parameters), there is $\psi\in\mathcal{L}_{\mathrm{oag}}(vK)$ which defines $D$.
The following lemma constitutes the core of the arguments of Jahnke and Koenigsmann, and will be used repeatedly throughout our proofs as well. It relies on a classical fact about divisible ordered abelian groups, which follows from quantifier elimination.
Fact 3.1.3 ().
Let $\Gamma$ be a divisible ordered abelian group. Then $\Gamma$ has no proper, $\mathcal{L}_{\mathrm{oag}}$-definable (even with parameters) non-trivial convex subgroup.
Lemma 3.1.4 ().
Suppose $(K,v)$ is a tame valued field of positive residue characteristic, with $vK$ divisible. If $w$ is a proper non-trivial coarsening of $v$, then $w$ cannot be $\mathcal{L}_{\mathrm{val}}$-definable. In particular, it cannot be $\mathcal{L}_{\mathrm{ring}}$-definable.
Proof.
Suppose $w$ is $\mathcal{L}_{\mathrm{val}}$-definable (possibly with parameters); then the corresponding proper non-trivial convex subgroup
$\Delta=\{x\in vK\mid\exists y(v(y)=x\land w(y)=0)\},$
is $\mathcal{L}_{{K,\Gamma,k}}$-definable (possibly with parameters) as a subset of $vK$. By Fact 3.1.2, $\Delta$ is $\mathcal{L}_{\mathrm{oag}}(vK)$-definable. This is a contradiction, as divisible ordered abelian groups have no definable proper non-trivial convex subgroups. ∎
The proof of the following theorem is adapted directly from [JK17, Corollary 6.1].
Theorem 3.1.5 ().
Let $(K,v_{K})$ be such that $Kv_{K}$ is perfect of positive characteristic. Assume that $K$ admits a definable non-trivial henselian valuation. Then, at least one of the following holds:
  1. (1)
    $Kv_{K}$ is separably closed,
  2. (2)
    $Kv_{K}$ is not t-henselian,
  3. (3)
    there is $L\succeq Kv_{K}$ such that $v_{L}L$ is not divisible,
  4. (4)
    $v_{K}K$ is not divisible,
  5. (5)
    $(K,v_{K})$ is not defectless,
  6. (6)
    there is $L\succeq Kv_{K}$ such that $(L,v_{L})$ is not defectless.
Proof.
We assume
$\neg 1\land\neg 2\land\neg 3\land\neg 4\land\neg 5\land\neg 6,$
and want to show that $K$ then cannot admit a non-trivial definable henselian valuation. Note that $\neg 1$ means that $H_{2}(K)=\varnothing$. Thus, all henselian valuations on $K$ are linearly ordered and coarser than $v_{K}$. Now, let ${(M,v)\succeq(K,v_{K})}$ be an $\aleph_{1}$-saturated extension in $\mathcal{L}_{\mathrm{val}}$. Then, ${Mv\succeq Kv_{K}}$ is an $\aleph_{1}$-saturated elementary extension of a t-henselian field (by $\neg 2$), hence it is henselian by Fact 2.5.1. Since ${L:=Mv\equiv Kv_{K}}$ is henselian and not separably closed (by $\neg 1$), $v_{L}$ is non-trivial and in fact, by Remark 2.3.1, we have the following diagram (where the arrows are intended as places),
[…]
from which we can deduce that $v_{M}=v_{L}\circ v$ is a proper refinement of $v$. Now, since $v_{L}L$ is divisible (because of $\neg 3$), so is $v_{M}M$. Indeed,
$vM={v_{M}M}/{v_{L}L},$
and we know $vM$ to be divisible, as $v_{K}K$ is (because of $\neg 4$).
Now, by assumption $Kv_{K}$ is a perfect field of positive characteristic and thus so is $L=Mv\succeq Kv_{K}$. It follows that also the residue field $Lv_{L}=Mv_{M}$ of the valuation $v_{L}$ on $L$ is perfect. By assumptions $\neg 5$ and $\neg 6$, $v$ and $v_{L}$ are defectless. Thus $v_{M}$ is defectless (as the composition of defectless valuations is defectless, see [AJ, Lemma 2.9]), and $(M,v_{M})$ is a tame valued field.
Assume now that there is a non-trivial definable henselian valuation $u$ on $K$, with valuation ring defined by the $\mathcal{L}_{\mathrm{ring}}(K)$-formula $\psi(x)$. Because of $\neg 1$, $u$ is a (not necessarily proper) coarsening of $v_{K}$. Now, via the elementary embedding, $\psi(x)$ defines a valuation $u^{*}$ on $M$, which is a coarsening of $v$ and hence a proper coarsening of $v_{M}$. This is a contradiction to Lemma 3.1.4. ∎
3.2. Failure of divisibility as a source of definability
We now seek to extend the analysis of Jahnke and Koenigsmann of the case when $v_{K}K$ is itself divisible, but there is an elementary extension $L$ of $Kv_{K}$ which is henselian with non-divisible value group $v_{L}L$.
Proposition 3.2.1 ().
Let $K$ be a henselian field, and let $v_{K}$ be its canonical henselian valuation. If
  1. (1)
    there is $L\equiv Kv_{K}$ such that $v_{L}L$ is non-divisible, and
  2. (2)
    $v_{K}K$ is divisible,
then $K$ admits an $\varnothing$-definable non-trivial henselian valuation.
Proof.
Let $L\equiv Kv_{K}$ be such that $v_{L}L$ is non-divisible. First, we notice that $Kv_{K}$ is not separably closed. Otherwise, $L$ would also be separably closed, but separably closed valued fields always have divisible a value group, and $v_{L}L$ is not divisible by assumption. Thus $v_{K}\in H_{1}(K)$ and all henselian valuations on $K$ are coarsenings of $v_{K}$. If $Kv_{K}$ would admit a non-trivial henselian valuation, then composing it with $v_{K}$ would give rise to a proper henselian refinement of $v_{K}$, which is not possible. Hence $Kv_{K}$ does not admit any non-trivial henselian valuation. We claim that $v_{L}\in H_{1}(L)$. Assume for a contradiction that it is not. Then by Theorem 2.6.1 there is an $\varnothing$-definable non-trivial henselian valuation on $L$. Thus there is also an $\varnothing$-definable non-trivial henselian valuation $u$ on $Kv_{K}$ by elementary equivalence, contradicting the considerations in the previous paragraph. By the Keisler-Shelah isomorphism theorem [Mar06, Theorem 2.5.36], there are an index set $I$ and an ultrafilter $\mathcal{U}$ on $I$ such that $L^{\mathcal{U}}\cong(Kv_{K})^{\mathcal{U}}$. Take the ultrapower $(L^{\mathcal{U}},v_{L}^{\mathcal{U}}):=(L,v_{L})^{\mathcal{U}}$ as an $\mathcal{L}_{\mathrm{val}}$-structure. Then, $v_{L}^{\mathcal{U}}L^{\mathcal{U}}\equiv v_{L}L$ is not divisible. Since $v_{L}\in H_{1}(L)$, then $v_{L}^{\mathcal{U}}\in H_{1}(L^{\mathcal{U}})$ and thus $v_{L}^{\mathcal{U}}$ coarsens the canonical henselian valuation $v_{L^{\mathcal{U}}}$ on $L^{\mathcal{U}}$. In particular, also $v_{L^{\mathcal{U}}}L^{\mathcal{U}}$ cannot be divisible, say not $q$-divisible for some prime $q$. Moreover, if we let $(K^{\mathcal{U}},v_{K}^{\mathcal{U}}):=(K,v_{K})^{\mathcal{U}}$ as an $\mathcal{L}_{\mathrm{val}}$-structure, then $K^{\mathcal{U}}v_{K}^{\mathcal{U}}=L^{\mathcal{U}}$ and $v_{K}^{\mathcal{U}}K^{\mathcal{U}}\equiv v_{K}K$ is divisible. Now,
  • $\circ$
    $K^{\mathcal{U}}$ is henselian and $v_{K^{\mathcal{U}}}=v_{L^{\mathcal{U}}}\circ v_{K}^{\mathcal{U}}$ (Remark 2.3.1),
  • $\circ$
    $v_{L^{\mathcal{U}}}L^{\mathcal{U}}\subseteq v_{K^{\mathcal{U}}}K^{\mathcal{U}}$ is a convex subgroup and thus also $v_{K^{\mathcal{U}}}K^{\mathcal{U}}$ is not $q$-divisible, and
  • $\circ$
    $v_{K^{\mathcal{U}}}K^{\mathcal{U}}/v_{L^{\mathcal{U}}}L^{\mathcal{U}}\cong v_{K}^{\mathcal{U}}K^{\mathcal{U}}$ is divisible, hence $v_{K^{\mathcal{U}}}K^{\mathcal{U}}$ is not $q$-antiregular.
We can then apply Theorem 2.6.1 to $(K^{\mathcal{U}},v_{K^{\mathcal{U}}})$ (note the different valuation!) and find an $\varnothing$-definable non-trivial henselian valuation on $K^{\mathcal{U}}$. Since ${K\preceq K^{\mathcal{U}}}$, the same $\mathcal{L}_{\mathrm{ring}}$-formula defines a non-trivial henselian valuation ring on $K$. ∎
4. Defect
Defect – i.e. the quantity $d$ in equation ($\star$2.4) – measures how much of the finite extension of a henselian valued field $(K,v)$ doesn’t correspond to extension of its value group $vK$ and of its residue field $Kv$. Originally introduced by Ostrowski (see [Roq02] for a thorough summary of his work), it was then extensively studied by Franz-Viktor Kuhlmann in his thesis [Kuh90] and his subsequent work. While in classical number theoretic situations, defect is trivial, it is a crucial invariant in the model theory of valued fields, in a certain sense limiting the Ax-Kochen/Ershov philosophy. In residue characteristic zero, every extension is defectless; this is why, for example, the classical Ax-Kochen/Ershov theorem can be proven for all henselian equicharacteric zero valued fields.
Unlike the equicharacteristic zero situation, valued fields of positive residue characteristic showcase wildly different behaviour. In particular, the fundamental equality can fail, i.e. there might be an extension with $d>1$. This is usually a serious obstruction to most efforts in understanding complete theories of valued fields. Throughout this section, however, we will deploy such extensions to define henselian valuations.
4.1. A toy situation
Suppose that $v\in H_{1}(K)$ and there exists an $\mathcal{L}_{\mathrm{ring}}$-definable subset $D$ of $K$ with
$v(D)=\{\gamma\in vK\mid\gamma>H\},$
where $H$ is a proper convex subgroup of $vK$.
{tikzpicture}
We will prove that whenever $(K,v)$ is as above, then there is a definable coarsening of $v$ which corresponds to $H$, see 2.2. This is a simplified version of the situation that will show up later when working with independent defect extensions. All the main ideas are already contained in this proof, with the advantage that we can avoid the technical subtleties of working modulo an interpretation of a finite field extension, which make the proof of Theorem 4.3.1 cumbersome.
Remark 4.1.1 ().
We will repeatedly use Beth’s definability theorem ([Hod97, Theorem 5.5.4]). In our case, this means that in order to show that a certain subset $D\subseteq K^{n}$ is $\mathcal{L}_{\mathrm{ring}}(\overline{c})$-definable on $K$, it is enough to work in the augmented language ${\mathcal{L}=\mathcal{L}_{\mathrm{ring}}(\overline{c})\cup\{P(X_{1},\dots X_{n})\}}$, and show that that whenever we take two structures ${(L,\overline{c}^{\prime},D_{1}),(L,\overline{c}^{\prime},D_{2})\equiv_{\mathcal{L}}(K,\overline{c},D)}$, we have $D_{1}=D_{2}$. In particular, in the proof of Lemma 4.1.2, we will consider valuation rings and thus unary predicates.
Lemma 4.1.2 ().
Let $(K,v)$ be a valued field such that $v\in H_{1}(K)$. Let $D\subseteq K$ be an $\mathcal{L}_{\mathrm{ring}}$-definable subset such that, for some proper convex subgroup $H\subseteq vK$ (possibly equal to the trivial subgroup),
$\{\alpha\in vK\mid\alpha>H\}=\{v(d)\mid d\in D\}.$
Let $v_{H}$ be the coarsening of $v$ corresponding to $H$ and denote by $\mathcal{O}_{H}$ its valuation ring. Then $v_{H}$ is definable.
Proof.
We will use Beth’s definability theorem as explained in Remark 4.1.1. We work in $\mathcal{L}_{\mathrm{val}}$: let $L$ and $\mathcal{O}_{1},\mathcal{O}_{2}\subseteq L$ be such that ${(L,\mathcal{O}_{1}),(L,\mathcal{O}_{2})\equiv(K,\mathcal{O}_{H})}$. In particular, both $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$ are henselian valuation rings on $L$, corresponding to valuations $v_{1}$ and $v_{2}$. We want to show that $\mathcal{O}_{1}=\mathcal{O}_{2}$. Note that, since $v_{H}\in H_{1}(K)$, the same holds for $v_{1}$ and $v_{2}$, and thus $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$ are comparable. Without loss of generality, assume $\mathcal{O}_{1}\subseteq\mathcal{O}_{2}$. This implies that $\mathcal{O}_{1}^{\times}\subseteq\mathcal{O}_{2}^{\times}$. We note that, if $D=\psi(K)$ for some $\mathcal{L}_{\mathrm{ring}}$-formula $\psi$ (possibly with parameters), then
$(K,\mathcal{O}_{H})\vDash\forall x\left(x\notin\mathcal{O}_{H}\iff\exists u\exists y\left(u\in\mathcal{O}_{H}^{\times}\land\psi(y)\land\left(x=u\frac{1}{y}\right)\right)\right).$
The same $\mathcal{L}_{\mathrm{val}}$-sentence is then true in both $(L,\mathcal{O}_{1})$ and $(L,\mathcal{O}_{2})$.
Now, towards proving $\mathcal{O}_{1}=\mathcal{O}_{2}$, assume for a contradiction that there is $x\in\mathcal{O}_{2}$ but $x\notin\mathcal{O}_{1}$: then, there exist $u\in\mathcal{O}_{1}^{\times}\subseteq\mathcal{O}_{2}^{\times}$ and $d\in\psi(L)$ such that $x=u\frac{1}{d}$. In particular, then, by the $\mathcal{L}_{\mathrm{val}}$-sentence above, $x\notin\mathcal{O}_{2}$, a contradiction. It follows that $\mathcal{O}_{1}=\mathcal{O}_{2}$.
We can apply Beth’s definability theorem and get that $D=\varphi(K)$ for some $\mathcal{L}_{\mathrm{ring}}$-formula $\varphi(x)$. ∎
Remark 4.1.3 ().
Beth’s definability theorem grants some control over parameters; namely, if $D$ was $\varnothing$-definable to begin with, then so is $\mathcal{O}_{H}$. Moreover, the same argument works if we replace $\mathcal{L}_{\mathrm{ring}}$ with some expansion $\mathcal{L}$; we then get that $v_{H}$ is $\mathcal{L}$-definable.
4.2. Independent defect
Independent defect will be the central tool of our definability proof; the set $\Sigma_{L}$ which will be defined in a moment, while not $\mathcal{L}_{\mathrm{ring}}$-definable per se, will contain enough information to $\mathcal{L}_{\mathrm{ring}}$-define a henselian valuation ring modulo the interpretation of a finite field extension.
Note that every equicharacteristic 00 valued field is defectless, so from now on for the rest of this section we assume that our valued fields have residue characteristic $p>0$.
Definition 4.2.1 ([KR23, Introduction]).
Let $(K,v)\subseteq(L,v)$ be a Galois extension of degree $p$ with defect. For any $\sigma\in\operatorname{Gal}(L|K)\setminus\{\mathrm{id}\}$, we let
$\Sigma_{\sigma}\coloneqq\left\{v\left(\frac{\sigma f-f}{f}\right)\mathrel{\Big{|}}f\in L^{\times}\right\}.$
Fact 4.2.2 ([KR23, Theorems 3.4 and 3.5]).
$\Sigma_{\sigma}$ does not depend on the choice of $\sigma$, and is a final segment of ${vK}_{\infty}$. We denote it by $\Sigma_{L}$.
Definition 4.2.3 ([KR23, Introduction]).
The extension $(K,v)\subseteq(L,v)$ has independent defect if there is a proper convex subgroup $H\subseteq vK$ such that $vK/H$ does not have a minimum positive element, and further
$\Sigma_{L}=\{\alpha\in{vK}_{\infty}\mid\alpha>H\}.$
We say that the extension has dependent defect otherwise.
Definition 4.2.4 ([KR23, Introduction]).
Let $(K,v)$ be a henselian valued field of positive residue characteristic $p$. If $\operatorname{char}(K)=0$, then let $K^{\prime}\coloneqq K(\zeta_{p})$ where $\zeta_{p}$ is a primitive $p$th root of unity. Otherwise let $K^{\prime}\coloneqq K$. We denote the unique extension of $v$ to $K^{\prime}$ also by $v$. We say $(K,v)$ is an independent defect field if all Galois defect extensions of $(K^{\prime},v)$ of degree $p$ have independent defect.
We now prove a result on composition of independent defect valuations with defectless valuations, which will be quite useful later. We couple it with an open question, which seemed natural enough.
Lemma 4.2.5 ().
Let $(K,v)$ be a henselian valued field with $v=\bar{v}\circ w$, such that $(K,w)$ is defectless and $(Kw,\bar{v})$ is a perfect independent defect field. Then $(K,v)$ is an independent defect field.
Proof.
Assume that $v=\overline{v}\circ w$, with $w$ defectless and $(Kw,\overline{v})$ an independent defect field. If $(K,v)$ has mixed characteristic $(0,p)$ we assume without loss of generality that $K$ contains a primitive $p$-th root of unity.
[…]
contradicting $[Lw:Kw](wL:wK)=p$.
Using this and $vL=vK$, we get
[…]
$\Sigma_{Lw}\coloneqq\left\{\bar{v}\left(\frac{f-\tau(f)}{f}\right)\mathrel{\Big{|}}f\in(Lw)^{\times}\right\}=\{\gamma\in{\bar{v}(Kw)}_{\infty}\mid\gamma>H\},$
[…]
$\Sigma_{L}\coloneqq\left\{v\left(\frac{f-\sigma(f)}{f}\right)\mathrel{\Big{|}}f\in L^{\times}\right\}=\{\gamma\in{vK}_{\infty}\mid\gamma>H\},$
where $\sigma\in\operatorname{Gal}(L|K)$ is such that $\langle\sigma\rangle=\operatorname{Gal}(L|K)$. This will prove that the defect in the extension $(L,v)\supseteq(K,v)$ is independent.
As $\Sigma_{L}$ is a final segment in ${vK}_{\infty}$ (by Fact 4.2.2), it is enough to show that $\Sigma_{L}\cap{\bar{v}(Kw)}_{\infty}=\Sigma_{Lw}$. Note that since $(K,w)$ is henselian, $\mathcal{O}_{(L,w)}$ is fixed setwise by the action of $\operatorname{Gal}(L|K)$. Thus by [EP05, Proposition 3.2.16(3)], $\phi$ induces an automorphism $\bar{\phi}\in\operatorname{Gal}(Lw|Kw)$ given by
$\bar{\phi}(\operatorname{res}_{w}(f))\coloneqq\operatorname{res}_{w}(\phi(f)),\quad f\in\mathcal{O}_{L}.$
We argue that the map
$\left\{\begin{array}[]{rcl}\operatorname{Gal}(L|K)&\mathchoice{\longrightarrow}{\rightarrow}{\rightarrow}{\rightarrow}&\operatorname{Gal}(Lw|Kw)\\ \phi&\mapsto&\bar{\phi}\end{array}\right.$
is a surjective group homomorphism (cf. the proof of [EP05, Lemma 5.2.6]): indeed, we write $Lw=Kw(\operatorname{res}_{w}(a))$, and we take any $\rho\in\operatorname{Gal}(Lw|Kw)$. We let $a_{1},\dots a_{p}$ be the $\operatorname{Gal}(L|K)$-conjugates of $a$.
For some $i\leq p$, ${\operatorname{res}_{w}(a_{i})=\rho(\operatorname{res}_{w}(a))}$, so there is $\sigma\in\operatorname{Gal}(L|K)$ such that
${\operatorname{res}_{w}(\sigma(a))=\bar{\sigma}(\operatorname{res}_{w}(a))=\rho(\operatorname{res}_{w}(a))}.$
Thus $\bar{\sigma}$ and $\rho$ coincide on $Lw$.
Now, since $(\phi\mapsto\bar{\phi})$ is surjective between finite groups with the same order, it is an isomorphism. It follows that it maps generators of $\operatorname{Gal}(L|K)$ to generators of $\operatorname{Gal}(Lw|Kw)$.
We can now compare the sets $\Sigma_{L}$ and $\Sigma_{Lw}$.
Step 1: $\Sigma_{Lw}\subseteq\Sigma_{L}\cap{\bar{v}(Kw)}_{\infty}$. Fix $\sigma\in\operatorname{Gal}(L|K)$ such that $\langle\sigma\rangle=\operatorname{Gal}(L|K)$. Then, since the set $\Sigma_{Lw}$ does not depend on the choice of the generator, we can write
$\Sigma_{Lw}$ =$=\left\{\bar{v}\left(\frac{f-\bar{\sigma}(f)}{f}\right)\mathrel{\Big{|}}f\in(Lw)^{\times}\right\}$
=$=\left\{\bar{v}\left(\frac{\operatorname{res}_{w}(f)-\bar{\sigma}(\operatorname{res}_{w}(f))}{\operatorname{res}_{w}(f)}\right)\mathrel{\Big{|}}f\in\mathcal{O}_{(L,w)}^{\times}\right\}$
=$=\left\{\bar{v}\left(\operatorname{res}_{w}\left(\frac{f-\sigma(f)}{f}\right)\right)\mathrel{\Big{|}}f\in\mathcal{O}_{(L,w)}^{\times}\right\}$
=$=\left\{v\left(\frac{f-\sigma(f)}{f}\right)\mathrel{\Big{|}}f\in\mathcal{O}_{(L,w)}^{\times}\right\}\cap{\bar{v}(Kw)}_{\infty}\subseteq\Sigma_{L}\cap{\bar{v}(Kw)}_{\infty},$
since $\bar{v}(\operatorname{res}_{w}(x))=v(x)$ for $x\in L$ such that $v(x)\in\bar{v}(Kw)$.
Step 2: $\Sigma_{Lw}\supseteq\Sigma_{L}\cap{\bar{v}(Kw)}_{\infty}$. Let $v\Big{(}\frac{f-\sigma(f)}{f}\Big{)}\in\Sigma_{L}\cap\bar{v}(Kw)$, i.e. $f\in L^{\times}$, then because $wL=wK$, there is $g\in K^{\times}$ with $w(f)=w(g)$ and so for $h\coloneqq\frac{f}{g}\in\mathcal{O}_{(L,w)}^{\times}$ we have
$\frac{f-\sigma(f)}{f}=\frac{h-\sigma(h)}{h},$
so
$v\left(\frac{f-\sigma(f)}{f}\right)=v\left(\frac{h-\sigma(h)}{h}\right)\in\Sigma_{Lw}.\qed$
Question 4.2.6 ().
Let $(K,v)$ be a henselian valued field with $v=\bar{v}\circ w$, such that both $(K,w)$ and $(Kw,\bar{v})$ are independent defect fields. Must then $(K,v)$ be an independent defect field?
4.3. Defining valuations from independent defect
We now have all the tools needed to build a definable henselian valuation out of a (independent) defect extension. This subsection is structured in several steps:
  • $\circ$
    Theorem 4.3.1 contains the core of the argument, showing how to deploy an independent defect extension (and its associated convex subgroup) to build a definable valuation.
  • $\circ$
    Corollary 4.3.3 shows how to go from our assumptions in the Main Theorem to the situation where the right defect extension actually appears.
  • $\circ$
    Corollaries 4.3.5 and 4.3.6 then specialize to the mixed characteristic situation, where some extra care is needed.
Finally, in Subsection 4.4, we will deal with a scenario that mirrors the one of Proposition 3.2.1, namely with the situation where $(K,v_{K})$ is defectless, but some elementary extension $L$ of $Kv_{K}$ is henselian and $v_{L}$ admits defect.
The following theorem should be thought of as a more elevated version of Lemma 4.1.2, where the $\mathcal{L}_{\mathrm{ring}}$-definable set $D$ now appears as a subset of the cartesian product $K^{p}$, which we identify along an interpretation with a degree $p$ Galois defect extension. For details on interpretations of finite field extensions, see for example [Cha09, Section 3.5].
Theorem 4.3.1 ().
Let $(K,v)$ be a henselian valued field with $\operatorname{char}{Kv}=p>0$ such that $(K,v)$ admits a Galois extension of degree $p$ with independent defect. If $\operatorname{char}(K)=0$, we additionally assume that $\zeta_{p}\in K$, where $\zeta_{p}$ is a primitive $p$-th root of unity. Then $K$ admits a non-trivial definable henselian valuation, coarsening $v$.
Proof.
Note that we can assume that $v\in H_{1}(K)$, since otherwise we could use [JK15, Theorem 3.10]. Let $K\subseteq L$ be the degree $p$ Galois extension with independent defect and take $\theta$ such that $L=K(\theta)$. Let $\sigma$ be a generator of the (cyclic) Galois group $\operatorname{Gal}(L|K)$, and let
$D\coloneqq\left\{\frac{\sigma(f)-f}{f}\mathrel{\Big{|}}f\in L^{\times}\right\}.$
As the extension has independent defect, there is a (possibly trivial) convex subgroup $H\subsetneq vK$ such that
$\Sigma_{L}=v(D)=\{\alpha\in{vK}_{\infty}\mid\alpha>H\}.$
Recall that $L=K(\theta)$ can be interpreted in $K$ via the $K$-linear isomorphism
$\left\{\begin{array}[]{rcl}K^{p}&\mathchoice{\longrightarrow}{\rightarrow}{\rightarrow}{\rightarrow}&L=K(\theta)\\ (a_{0},\ldots,a_{p-1})&\mapsto&\sum_{i=0}^{p-1}a_{i}\theta^{i}\end{array}\right.$
We will denote the above map with $f$. Note that the coefficients of the minimal polynomial of the generator $\theta$ over $K$ are needed as parameters to describe the multiplication. Now, the action of $\sigma$ is definable in $K$ (via the interpretation) using the coefficients of the change of basis matrix. If $\operatorname{char}(K)=p$ then the matrix has integer coefficients, so no additional parameter is needed. If $\operatorname{char}(K)=0$, then we need to use $\zeta_{p}$. We will denote the parameter tuple needed to define $D$ in $K$ by $\bar{c}$ (i.e. the coefficients of the minimal polynomial of the generator $\theta$ over $K$, and $\zeta_{p}$ if necessary). Thus, $f^{-1}(D)\subseteq K^{p}$ is definable in $K$, using $\overline{c}$ as parameters; choose an $\mathcal{L}_{\mathrm{ring}}(\overline{c})$-formula $\psi(\overline{X},\overline{c})$ so that $f^{-1}(D)=\psi(K^{p},\overline{c})$. Denote by $\mathcal{O}_{H}\subseteq L$ the coarsening of (the unique extension of) $v$ corresponding to $H$. We can see $\mathcal{O}_{H}$ as a subset $f^{-1}(\mathcal{O}_{H})$ of $K^{p}$.
Claim: $f^{-1}(\mathcal{O}_{H})\subseteq K^{p}$ is $\mathcal{L}_{\mathrm{ring}}(\overline{c})$-definable.
We will use Beth’s definability theorem. We work in ${\mathcal{L}\coloneqq\mathcal{L}_{\mathrm{ring}}(\overline{c})\cup\{P(\overline{X})\}}$, where $P(\overline{X})$ is a $p$-ary predicate. Under the bijection $f$ that establishes the interpretation, the predicate will represent the valuation subring $\mathcal{O}_{H}$ of $K(\theta)$. We now let $(M,\overline{c}^{\prime},P_{1}),(M,\overline{c}^{\prime},P_{2})\equiv_{\mathcal{L}}(K,\overline{c},f^{-1}(\mathcal{O}_{H}))$. Note that the interpretation given by $f$ gives rise to an $\mathcal{L}_{\mathrm{ring}}$-structure $N$ and a map $f^{\prime}$ which is establishing an interpretation of $N$ in $M$. Furthermore, because $(M,\overline{c}^{\prime},P_{1})$ and $(M,\overline{c}^{\prime},P_{2})$ are elementarily equivalent, we have that:
  • $\circ$
    $M\subseteq N$ are fields, and the polynomial with coefficients from $\overline{c}^{\prime}$ is of degree $p$ and irreducible over $M$; thus, the field extension $M\subseteq N$ is algebraic of degree $p$;
  • $\circ$
    $\mathcal{O}_{1}=f^{\prime}(P_{1})$ and $\mathcal{O}_{2}=f^{\prime}(P_{2})$ are non-trivial henselian valuation rings on $N$; we denote the corresponding valuations by $v_{1}$ and $v_{2}$, respectively;
  • $\circ$
    there is $\ell\geq 2$ such that there is a separable polynomial of degree $\ell$ over $Lv_{H}$ with no root in $Lv_{H}$; thus, there are separable polynomials of degree $\ell$ over $Nv_{1}$ and $Nv_{2}$ with no root, and thus $v_{1},v_{2}\in H_{1}(N)$,
  • $\circ$
    for $i=1,2$, $x\notin\mathcal{O}_{i}$ if and only if there are $u\in\mathcal{O}_{i}^{\times}$ and $y\in\psi(M^{p},\overline{c}^{\prime})$ such that $x=u\frac{1}{f^{\prime}(y)}$.
Since $\mathcal{O}_{1},\mathcal{O}_{2}\in H_{1}(N)$, they must be comparable. Assume, for example, that $\mathcal{O}_{1}\subseteq\mathcal{O}_{2}$. Towards a contradiction, assume now that there is $x\in\mathcal{O}_{2}$ such that $x\notin\mathcal{O}_{1}$. In particular, this means that there are $u\in\mathcal{O}_{1}^{\times}$ and $y\in\psi(M^{p},\overline{c}^{\prime})$ such that $x=u\frac{1}{f^{\prime}(y)}$. But since $\mathcal{O}_{1}^{\times}\subseteq\mathcal{O}_{2}^{\times}$, and thus $u\in\mathcal{O}_{2}^{\times}$, we get that $x\notin\mathcal{O}_{2}$, a contradiction. Then $\mathcal{O}_{1}=\mathcal{O}_{2}$, so $P_{1}=P_{2}$ and we can apply Beth’s definability. This means that there is an $\mathcal{L}_{\mathrm{ring}}$-formula $\psi(\overline{X},\overline{C})$, such that ${f^{-1}(\mathcal{O}_{H})=\psi(K^{p},\overline{c})}$. $\blacksquare_{\text{Claim}}$ Now, we can define the (non-trivial) restriction $\mathcal{O}_{H}\cap K$ as follows: given $z\in K$, we have that
$z\in\mathcal{O}_{H}\cap K\iff K\vDash\psi(z,\underbrace{0,\dots 0}_{(p-1)\text{-times}},\overline{c}).$
This exhibits a non-trivial definable henselian valuation ring on $K$. ∎
We can now use this theorem to define a non-trivial henselian valuation on certain valued fields with defect, but first we need to produce the Galois defect extensions required to exhibit independent defect.
Lemma 4.3.2 ().
Let $(K,v)$ be a perfect valued field of residue characteristic $p>0$. If $(K,v)$ is not defectless, then there is a finite extension $K\subseteq K^{\prime}$ such that $K^{\prime}$ admits a defect Galois extension of degree $p$.
Proof.
Take a defect extension $K\subseteq L$. In particular, $p^{n}\mid[L:K]$ for some $n\geq 1$. We let $N$ be the normal hull of $L$; note that $p^{n}\mid[N:K]$ and $K\subseteq N$ is Galois. Consider now $H\subseteq\operatorname{Gal}(N|K)$ to be the $p$-Sylow subgroup. Then $L^{\prime}\coloneqq N^{H}$ is such that $L^{\prime}\subseteq N$ is Galois and it is a tower $L^{\prime}=L_{0}\subseteq L_{1}\subseteq\cdots\subseteq L_{n}=N$ of normal degree $p$ extensions. Moreover, since $K\subseteq L^{\prime}$ has degree prime to $p$, it is defectless, hence $L^{\prime}\subseteq N$ is defect. In particular, there is some $\ell\leq n$ such that $L_{\ell}\subseteq L_{\ell+1}$ is a defect Galois extension of degree $p$, so we take $K^{\prime}=L_{\ell}$. ∎
Corollary 4.3.3 ().
Suppose $(K,v)$ is a henselian valued field with $\operatorname{char}{Kv}=p>0$ such that $(K,v)$ is not defectless, and such that either
  1. (1)
    $K$ is perfect, if $\operatorname{char}{K}=p>0$, or
  2. (2)
    $\mathcal{O}_{v}/p$ is semi-perfect, if $\operatorname{char}{K}=0$.
Then $K$ admits a non-trivial definable henselian valuation, coarsening $v$.
Proof.
We may assume that $v\in H_{1}(K)$; otherwise, we could use [JK15, Theorem 3.10] to define a non-trivial henselian valuation.
First, we note that from our hypothesis it follows that $(K,v_{K})$ also has defect. Moreover, in the case where $\operatorname{char}{K}=0$, we make the following two observations:
  • $\circ$
    since $K$ always admits a definable valuation if $v$ is finitely ramified, we may assume that it is infinitely ramified,
  • $\circ$
    the assumptions on $(K,v)$ imply that it is a roughly deeply ramified valued field in the sense of [KR23, p. 2696] (note that by [KR23, Lemma 4.1], we only need to check that $\mathcal{O}_{v_{K}}/p$ is semi-perfect, without looking at the completion of $K$). Thus, by [KR23, Theorem 1.8] any finite extension of it is also roughly deeply ramified.
If $\operatorname{char}{K}=0$, let $K_{0}=K(\zeta_{p})$, where $\zeta_{p}$ is some primitive $p$th root of unity. Otherwise, take $K_{0}=K$. By Lemma 4.3.2, we have a finite extension $K_{0}\subseteq K_{1}$ which admits a Galois defect extension of degree $p$. Since $K\subseteq K_{1}$ is a finite separable field extension, we find $b\in K_{1}$ such that $K_{1}=K(b)$. The field $K_{1}$ admits a Galois defect extension of degree $p$, i.e there is $c\in K_{1}$ which has no Artin-Schreier root in $K$ if $\operatorname{char}{K}=p$ (resp. no $p$th root, if $\operatorname{char}{K}=0$), and the extension $(K_{1},v)\subseteq(K_{1}(\theta),v)$, where $\theta^{p}-\theta=c$ (resp. $\theta^{p}=c$), is an immediate defect extension. Denote by $L=K_{1}(\theta)=K(b,\theta)$. Then, we have to distinguish two cases:
  • $\circ$
    if $\operatorname{char}{K}=p$, then $K_{1}$ is a perfect field,
  • $\circ$
    if $\operatorname{char}{K}=0$, then $K_{1}$ is a finite extension of a roughly deeply ramified field, and thus it is roughly deeply ramified itself.
In both cases, $K_{1}$ is an independent defect field (see [KR23, Theorem 1.10]), and thus the extension $(K_{1},v)\subseteq(L,v)$ is a degree $p$ defect extension with independent defect. By Theorem 4.3.1, then, $K_{1}$ admits a non-trivial definable henselian valuation $v_{H}$, with valuation ring defined by $\varphi(X,c)$. Now recall that $K_{1}=K(b)$ can be interpreted in $K$, using the coefficients of the minimal polynomial of $b$ as parameters. Thus, $\mathcal{O}_{H}$ is also definable in $K$ (i.e., as a subset of $K^{np}$), using as parameters the coefficients of the minimal polynomial of $b$ over $K$, and the coordinates of $c$ in the chosen $K$-basis of $K_{1}$. We write $\overline{\rho}$ for the tuple of parameters needed to define $\mathcal{O}_{H}$ in $K$. Let $\widetilde{\varphi}(\overline{X},\overline{Y})$ be an $\mathcal{L}_{\mathrm{ring}}$-formula defining $\mathcal{O}_{H}$ in $K$, i.e., $\widetilde{\varphi}(K,\overline{\rho})=\mathcal{O}_{H}\subseteq K^{np}$. The set $\mathcal{O}_{H}^{K}\coloneqq\mathcal{O}_{H}\cap K\subseteq K$ can be written as
$\mathcal{O}_{H}^{K}=\widetilde{\varphi}(K,{0,\ldots,0},\overline{\rho})\subseteq K$
with $np-1$ zeros. It is thus $\mathcal{L}_{\mathrm{ring}}$-definable in $K$ using the parameter tuple $\rho$. ∎
As we will later have to deal with definable valuations in some ultrapower of our field $K$, we immediately apply results of Anscombe and Jahnke from [AJ18] to get rid of parameters in our definition. Note that this is a crucial point: if we define a henselian valuation ring in some elementary extension $K\preceq K^{*}$, and we want to use the same $\mathcal{L}_{\mathrm{ring}}$-formula to define a henselian valuation ring on $K$, we have to deal with parameters somehow, as henselianity cannot be expressed with finitely many formulas.
Remark 4.3.4 ().
Note that if $(K,v)$ is a henselian valued field of mixed characteristic, then $v_{K}$ is also of mixed characteristic.
Corollary 4.3.5 ().
Suppose $(K,v)$ is a henselian valued field of mixed characteristic, such that $(K,v)$ is not defectless and $\mathcal{O}_{v}/p$ is semi-perfect. Then $K$ admits an $\varnothing$-definable non-trivial henselian valuation.
Proof.
We may assume that $v\in H_{1}(K)$: otherwise, we can use [JK15, Theorem 3.10] to find a $\varnothing$-definable non-trivial henselian valuation. Now, $v_{K}$ is also of mixed characteristic, thus [AJ18, Theorem 1.1.(B)] yields the existence of an $\varnothing$-definable non-trivial henselian valuation. ∎
Later, in the proof of Proposition 4.4.1, we will find ourselves in the situation where we deal with a defect refinement $v_{1}$ of a valuation $v_{K}^{\mathcal{U}}$, where $\mathcal{O}_{v_{K}^{\mathcal{U}}}/p$ is semi-perfect. A priori, the refinement itself will not be such that $\mathcal{O}_{v_{1}}/p$ is semi-perfect, thus allowing us to access independent defect – this is what happens in positive characteristic, where the base field stays perfect and so defect Galois extensions of degree $p$ are automatically of independent type. In mixed characteristic, we have to argue directly that the refinement is an independent defect field (using Lemma 4.2.5).
Lemma 4.3.6 ().
Suppose $(K,v)$ is a henselian valued field of mixed characteristic, such that $(K,v)$ is not defectless, and there is a coarsening $w$ of $v$ such that $(K,w)$ is defectless and $Kw$ is perfect of characteristic $p>0$. Then $K$ admits an $\varnothing$-definable non-trivial henselian valuation.
Proof.
We may assume that $v\in H_{1}(K)$: otherwise, we can use [JK15, Theorem 3.10] to find a $\varnothing$-definable non-trivial henselian valuation. Using Lemma 4.3.2, there is a finite extension $K(\zeta_{p})\subseteq K_{1}$ such that $(K_{1},v)$ admits a defect extension of degree $p$, and we find $b\in K_{1}$ such that $K_{1}=K(b)$. We now show that $(K_{1},v)$ is an independent defect field. For this we consider the decomposition of $(K_{1},v)$ with respect to the coarsening $(K_{1},w)$ and we denote the induced valuation on $K_{1}w$ by $\bar{v}$:
[…]
Note that $(K_{1},w)$ is defectless, because it is a finite extension of $(K,w)$ and $K_{1}w$ is perfect because it is a finite extension of $Kw$. From Lemma 4.2.5 it follows that also $(K_{1},v)$ is an independent defect field. By Theorem 4.3.1, then, $K_{1}$ admits a non-trivial definable henselian valuation $v_{H}$, with valuation ring defined by $\varphi(X,c)$. Since $K_{1}=K(b)$ can be interpreted in $K$ using the coefficients of the minimal polynomial of $b$ as parameters, $\mathcal{O}_{H}$ is also definable in $K$ (i.e., as a subset of $K^{np}$), using the coefficients of the minimal polynomial of $b$ over $K$, and the coordinates of $c$ in the chosen $K$-basis of $K_{1}$. We write $\overline{\rho}$ for the tuple of parameters needed to define $\mathcal{O}_{H}$ in $K$. Let $\widetilde{\varphi}(\overline{X},\overline{Y})$ be an $\mathcal{L}_{\mathrm{ring}}$-formula defining $\mathcal{O}_{H}$ in $K$, i.e., $\widetilde{\varphi}(K,\overline{\rho})=\mathcal{O}_{H}\subseteq K^{np}$. The set $\mathcal{O}_{H}^{K}\coloneqq\mathcal{O}_{H}\cap K\subseteq K$ can be written as
$\mathcal{O}_{H}^{K}=\widetilde{\varphi}(K,{0,\ldots,0},\overline{\rho})\subseteq K$
with $np-1$ zeros. It is thus $\mathcal{L}_{\mathrm{ring}}$-definable in $K$ using the parameter tuple $\rho$.
By our assumptions, $v_{K}$ is also of mixed characteristic, thus [AJ18, Theorem 1.1.(B)] yields the existence of an $\varnothing$-definable non-trivial henselian valuation. ∎
4.4. Defect in elementary extensions of $Kv_{K}$
The following proposition is a direct result of Corollary 4.3.5, while we will need some more work for positive characteristic.
Proposition 4.4.1 ().
Let $K$ be a henselian field such that $(K,v_{K})$ has mixed characteristic and $\mathcal{O}_{v_{K}}/p$ is semi-perfect. If
  1. (1)
    there is $L\equiv Kv_{K}$ such that $(L,v_{L})$ is not defectless, and
  2. (2)
    $(K,v_{K})$ is defectless,
then $K$ admits an $\varnothing$-definable non-trivial henselian valuation.
Proof.
Choose $L$ such that $L\equiv Kv_{K}$ and $(L,v_{L})$ is not defectless. By the Keisler-Shelah isomorphism theorem [Mar06, Theorem 2.5.36], take an index set $I$ and an ultrafilter $\mathcal{U}$ on $I$ such that $(Kv_{K})^{\mathcal{U}}\cong L^{\mathcal{U}}$. Consider the ultrapowers in $\mathcal{L}_{\mathrm{val}}$ $(K^{\mathcal{U}},v_{K}^{\mathcal{U}})\coloneqq(K,v_{K})^{\mathcal{U}}$ and $(L^{\mathcal{U}},v_{L}^{\mathcal{U}})\coloneqq(L,v_{L})^{\mathcal{U}}$. Since $v_{L}^{\mathcal{U}}$ is a valuation with defect, while $v_{K}^{\mathcal{U}}$ is defectless, the composition $v_{1}$ is a non-trivial henselian valuation on $K^{\mathcal{U}}$ with defect. The valuation $v_{1}$ is not defectless and of mixed characteristic, since $v_{K}^{\mathcal{U}}$ is, and its coarsening $v_{K}^{\mathcal{U}}$ is defectless and has perfect residue field $K^{\mathcal{U}}v_{K}^{\mathcal{U}}=L^{\mathcal{U}}$. Thus by Lemma 4.3.6, there is an $\varnothing$-definable valuation on $K^{\mathcal{U}}$. By elementary equivalence, we can find a $\varnothing$-definable henselian valuation on $K$. ∎
In the case of positive characteristic, eliminating the parameters from the definition of the henselian valuation can be tricky; indeed, a theorem along the lines of [AJ18, Theorem 1.1.(B)] will not be true in general in positive characteristic. However, in our setting (particularly because we can assume $v_{K}K$ to be divisible) we manage to get an analogous result, where we take a valuation defined in an elementary extension $K\preceq K^{*}$ (with parameters possibly in $K^{*}$) and produce a valuation defined with parameters in $K$, thus allowing us to “push it down”.
Fact 4.4.2 (cf. [Jah, Proposition 2.4]).
Let $(K,v)$ be a henselian valued field with non-separably closed nor real closed residue field. Then $v$ has an $\mathcal{L}_{\mathrm{ring}}$-definable refinement.
Proposition 4.4.3 ().
Let $K$ be a perfect henselian field of positive characteristic. If
  1. (1)
    $v_{K}K$ is divisible, and
  2. (2)
    there is $L\equiv Kv_{K}$ such that $(L,v_{L})$ is not defectless, and
  3. (3)
    $(K,v_{K})$ is defectless,
then $K$ admits a $\varnothing$-definable non-trivial henselian valuation.
Proof.
Note that we may assume that $Kv_{K}$ is not separably closed; otherwise, [JK15, Theorem 3.10] gives a definable non-trivial henselian valuation.
By the Keisler-Shelah isomorphism theorem [Mar06, Theorem 2.5.36], take an index set $I$ and an ultrafilter $\mathcal{U}$ on $I$ such that ${(Kv_{K})^{\mathcal{U}}\cong L^{\mathcal{U}}}$. Consider the ultrapowers in $\mathcal{L}_{\mathrm{val}}$ ${(K^{\mathcal{U}},v_{K}^{\mathcal{U}})\coloneqq(K,v_{K})^{\mathcal{U}}}$ and ${(L^{\mathcal{U}},v_{L}^{\mathcal{U}})\coloneqq(L,v_{L})^{\mathcal{U}}}$. Since $v_{L}^{\mathcal{U}}$ is a valuation with defect, while $v_{K}^{\mathcal{U}}$ is defectless, the composition $v_{1}$ is a non-trivial henselian valuation on $K^{\mathcal{U}}$ with defect. Thus, by Corollary 4.3.3, there is an $\mathcal{L}_{\mathrm{ring}}(K^{\mathcal{U}})$-definable non-trivial henselian valuation $v_{1}^{\prime}$ coarsening $v_{1}$, say given by $\varphi(K^{\mathcal{U}},\overline{\beta})$ for some parameters $\overline{\beta}\in(K^{\mathcal{U}})^{\ell}$, $\ell\geq 1$.
Since $Kv_{K}$ is not separably closed, Fact 4.4.2 yields an $\mathcal{L}_{\mathrm{ring}}(K)$-definable valuation $w$ on $K$ (not necessarily henselian!) such that $\mathcal{O}_{w}\subseteq\mathcal{O}_{v_{K}}$. Let $\psi(X,\overline{c})$ be the formula defining $w$. Denote by $w^{*}$ the valuation corresponding to $\psi(K^{\mathcal{U}},\overline{c})$. Note that since $(K,v_{K})\preceq(K^{\mathcal{U}},v_{K}^{\mathcal{U}})$, then we still have that $w^{*}$ is a proper refinement of $v_{K}^{\mathcal{U}}$.
As in the proof of [EP05, Theorem 2.3.4], given two valuation rings $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$, we will denote by
$\mathcal{O}_{1}\cdot\mathcal{O}_{2}\coloneqq\left\{\frac{x}{y}\mathrel{\Big{|}}x\in\mathcal{O}_{1},\;y\in\mathcal{O}_{1}\setminus\mathfrak{m}_{2}\right\}.$
Note that $\mathcal{O}_{1}\cdot\mathcal{O}_{2}$ contains both $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$ and is always a valuation ring. Indeed, it is the finest valuation ring containing both: if $\mathcal{O}_{1},\mathcal{O}_{2}\subseteq\mathcal{O}_{3}$, then for any $x\in\mathcal{O}_{1}$ and $y\in\mathcal{O}_{1}\setminus\mathfrak{m}_{2}$, one has that $y\notin\mathfrak{m}_{3}$, so $\frac{1}{y}\in\mathcal{O}_{3}$ and thus $\frac{x}{y}\in\mathcal{O}_{3}$. We now consider the $\mathcal{L}_{\mathrm{ring}}$-definable set
$X=\left\{\overline{b}\in(K^{\mathcal{U}})^{\ell}\mid\varphi(K^{\mathcal{U}},\overline{b})\;\text{is a valuation ring and}\;\mathcal{O}_{w^{*}}\cdot\varphi(K^{\mathcal{U}},\overline{b})\neq K^{\mathcal{U}}\right\}.$
For each $\overline{b}\in X$, we consider the $\mathcal{L}_{\mathrm{ring}}(\overline{b})$-definable valuation ring
$\mathcal{O}_{\overline{b}}\coloneqq\mathcal{O}_{w^{*}}\cdot\varphi(K^{\mathcal{U}},\overline{b}),$
with corresponding valuation $v_{\overline{b}}$. As all overrings of a given valuation are comparable, any $v_{\overline{b}}$ for $\overline{b}\in X$ is comparable with $\mathcal{O}_{v_{K}^{\mathcal{U}}}$.
Since $v_{K}^{\mathcal{U}}$ is a tame valuation with divisible value group, Lemma 3.1.4 implies that $v_{\overline{b}}$ cannot be a proper coarsening. In particular, then, $\mathcal{O}_{v_{\overline{b}}}\subseteq\mathcal{O}_{v_{K}^{\mathcal{U}}}$. It follows that the union
$\mathcal{O}\coloneqq\bigcup_{\overline{b}\in X}\mathcal{O}_{\overline{b}}$
is a non-trivial $\mathcal{L}_{\mathrm{ring}}(K)$-definable (note the parameters!) valuation ring on $K^{\mathcal{U}}$.
To show that $\mathcal{O}$ is henselian, we will show that $\mathcal{O}$ contains $\varphi(K^{\mathcal{U}},\overline{\beta})$. It is enough to show that ${\mathcal{O}_{\overline{\beta}}\neq K^{\mathcal{U}}}$. However, since $\varphi(K^{\mathcal{U}},\overline{\beta})$ is comparable with $v_{K}^{\mathcal{U}}$, $\mathcal{O}_{\overline{\beta}}$ is either equal to $\varphi(K^{\mathcal{U}},\overline{\beta})$ (if $\varphi(K^{\mathcal{U}},\overline{\beta})$ coarsens $v_{K}^{\mathcal{U}}$), or to a refinement of $v_{K}^{\mathcal{U}}$. Either way, it is non-trivial.
We have just showed that $\mathcal{O}$ is a coarsening of $\varphi(K^{\mathcal{U}},\overline{\beta})$, thus it is henselian. In particular, the same formula defines a non-trivial henselian valuation on $K$, as needed. ∎
5. Proof of the main theorem
Main theorem .
Let $K$ be perfect, not separably closed, and henselian. If ${\operatorname{char}{K}=0}$ and ${\operatorname{char}{Kv_{K}}=p>0}$, assume that $\mathcal{O}_{v_{K}}/p$ is semi-perfect. Then $K$ admits a definable non-trivial henselian valuation if and only if at least one of the following conditions hold:
  1. (1)
    $Kv_{K}$ is separably closed,
  2. (2)
    $Kv_{K}$ is not t-henselian,
  3. (3)
    there is $L\succeq Kv_{K}$ such that $v_{L}L$ is not divisible,
  4. (4)
    $v_{K}K$ is not divisible,
  5. (5)
    $(K,v_{K})$ is not defectless,
  6. (6)
    there is $L\succeq Kv_{K}$ such that $(L,v_{L})$ is not defectless.
Proof.
If $\operatorname{char}{Kv_{K}}=0$, conditions $4$ and $5$ are trivial, and so the statement is exactly Theorem 1.0.1.
Assume $\operatorname{char}{K}v_{K}=p>0$ and note that $Kv_{K}$ is perfect: in mixed characteristic, $Kv_{K}$ is perfect since $\mathcal{O}_{v_{K}}/p$ is semi-perfect; in positive characteristic, $Kv_{K}$ is perfect since $K$ is. Thus, one direction is Theorem 3.1.5.
In the other direction, we do a case distinction. First, assume $1\vee 2\vee 4$, then it follows from Theorem 2.6.1 that $K$ admits a definable non-trivial henselian valuation.
If $3\wedge\neg 4$ then Proposition 3.2.1 yields a existance of definable non-trivial henselian valuation.
Suppose now that $5$ holds, i.e. $(K,v_{K})$ is not defectless. Then we get a non-trivial definable henselian valuation from Corollary 4.3.3.
Finally, assume that $6$ holds, and since we have already established the case where $1\vee 2\vee 3\vee 4\vee 5$ holds, we can also assume $\neg 1\land\neg 2\land\neg 3\land\neg 4\land\neg 5$. Then by Proposition 4.4.1 (in mixed characteristic) and Proposition 4.4.3 (in positive characteristic), there is an $\varnothing$-definable non-trivial henselian valuation on $K$. ∎